# Non-toric brane webs, Calabi–Yau 3-folds, and 5d SCFTs

### v1

@@ 4.6. Generic consistent webs and fully pushed 7-branes
Let $W$ be a generic consistent web of 5-branes with 7-branes. By Lemma 4.16, since $W$ is consistent, there exists a modified web of 5-branes with 7-branes $W^{\prime}$, obtained from $W$ by a finite sequence of pushing operation of 7-branes, such that the corresponding line bundle $\mathcal{L}^{\prime}$ on the degeneration of log Calabi–Yau surfaces $(\mathcal{Y}^{\prime},\mathcal{D}^{\prime},\mathcal{L}^{\prime})\rightarrow\mathbb{C}$ is nef. By the base point free theorem, there exists a morphism $c:(\mathcal{Y}^{\prime},\mathcal{D}^{\prime},\mathcal{L}^{\prime})\rightarrow((\mathcal{Y}^{\prime})^{\mathrm{pol}},(\mathcal{D}^{\prime})^{\mathrm{pol}},(\mathcal{L}^{\prime})^{\mathrm{pol}})$ contracting all curves $C^{\prime}$ in $\mathcal{Y}^{\prime}$ such that $\mathcal{L}^{\prime}\cdot C^{\prime}=0$, and so that the line bundle $(\mathcal{L}^{\prime})^{\mathrm{pol}}$ is ample. The polyhedral decomposition defined by the web $W^{\prime}$ is the dual intersection complex of the central fiber $((\mathcal{Y}^{\prime})^{\mathrm{pol}}_{0},(\mathcal{D}^{\prime})^{\mathrm{pol}}_{0},(\mathcal{L}^{\prime})^{\mathrm{pol}}_{0})$. Since $W^{\prime}$ is generic, the irreducible components of $((\mathcal{Y}^{\prime})^{\mathrm{pol}}_{0},(\mathcal{D}^{\prime})^{\mathrm{pol}}_{0},(\mathcal{L}^{\prime})^{\mathrm{pol}}_{0})$ are described by Lemma 4.27, and it follows that the 7-branes are disjoint from the web $W^{\prime}$, and that the vertices of $W^{\prime}$ are 3-valent of multiplicity one. To end the proof that the 7-branes are fully pushed, it remains to check condition ii) in Definition 4.28. If this condition were not satisfied, there would exist a $(p_{i},q_{i})$ 7-brane $x_{i}$ such that the half-line $x_{i}+\mathbb{R}_{<0}(p_{i},q_{i})$ does not intersect $W^{\prime}$. Then, there would exist an interior $(-1)$-curve $E$ in a corner blow-up of $Y$ whose tropicalization is $x_{i}+\mathbb{R}_{<0}(p_{i},q_{i})$, and such that $L\cdot E=0$. But since $W$ is a consistent web, we have $L\cdot E>0$ by Definition 4.5 ii), which is a contradiction.
@@ 5.1. Consistent decorated toric polygons
Let $(\overline{P},(L_{i})_{1\leq i\leq Q})$ be a decorated toric polygon with a regular polyhedral decomposition $\overline{\mathcal{P}}$. We say that a web of 5-branes with 7-branes $W$ is dual to $(\overline{P},(L_{i})_{1\leq i\leq Q},\overline{\mathcal{P}})$ if there exists a piecewise-linear function $\varphi$ on $\overline{P}$ with domains of linearity $\overline{\mathcal{P}}$, such that $W$ is obtained from the web of 5-brane $\overline{W}$ which is dual to $(\overline{\mathcal{P}},\varphi)$ by adding for every $1\leq i\leq Q$ a 7-brane ending the leg of $\overline{W}$ corresponding to $L_{i}$. For any decorated toric polygon with a regular polyhedral decomposition, there is such a dual web associated to it, which is unique up to deformation.
Recall from §3.3 that there is a log Calabi surface with line bundle $(Y,D,L)$ associated to any web of $5$-branes with $7$-branes. In the following subsections, when we say the log Calabi surface with line bundle associated to a decorated toric polygon with a regular polyhedral decomposition, we mean the data of $(Y,D,L)$ associated to the dual web.
@@
    For every $1\leq i\leq Q$, a set of faces $\mathcal{F}_{\Psi,i}\subseteq\overline{\mathcal{P}}^{[2]}$. We refer to a face in $\mathcal{F}_{\Psi,i}$ as a “$(\Psi,i)$-colored face”, and require that $\mathcal{F}_{\Psi,i}\cap\mathcal{F}_{\Psi,j}=\emptyset$, if $i\neq j$.
@@
    For every $1\leq i\leq Q$, an element $A_{\Psi,i}\in\overline{\mathcal{P}}^{[0]}\cup\overline{\mathcal{P}}^{[1]}$, referred to as the “$(\Psi,i)$-active cell”.
@@
    ${\mathcal{F}_{\Psi_{\mathrm{in}},i}}=\emptyset$, and;
@@
    for all $11\leq j\leq k-$ 1, $\sigma_{j}$ is an elementary trapezoid, with two parallel edges $e_{in,j}$ and $e_{out,j}$, which were once active cells, and with two other edges $e_{l,j}$ and $e_{r,j}$ such that $e_{l,j}\sim_{\Psi}e_{r,j}$.
@@
Moreover, we have $|e_{in,1}|=|L_{i}|$, and, for every $11\leq j\leq k-$ 1, we have $|e_{out,j}|=|e_{in,j+1}|$.
@@ 5.2. Symington polygons
As in [37, Lemma 5.6], the integral affine structure with singularities on $P$ can be perturbed to obtain an integral affine manifold $\widetilde{P}$ with only focus-focus singularities. Then, as explained in [93, §5.4], one can view $\widetilde{P}$ as the base of an almost toric fibration on $(Y^{\mathrm{pol}},D^{\mathrm{pol}},L^{\mathrm{pol}})$, that is, a fibration whose general fibers are tori $T^{2}$, and singular fibers are pinched tori. Moreover, the fibers are Lagrangian submanifolds with respect to a symplectic form $\omega$ on $Y$ with class in $c_{1}(L^{\mathrm{pol}})\in H^{2}(Y^{\mathrm{pol}},\mathbb{R})$. This fibration is obtained by surgeries introducing singular fibers in the toric momentum map fibration $(\overline{Y},\overline{D},\overline{L})\rightarrow\overline{P}$ on the polarized toric surface $(\overline{Y},\overline{D},\overline{L})$ with momentum polytope $\overline{P}$. An alternative way to understand such almost toric fibrations is to consider the degeneration $((\mathcal{Y}^{\prime})^{\mathrm{pol}},(\mathcal{D}^{\prime})^{\mathrm{pol}},(\mathcal{L}^{\prime})^{\mathrm{pol}})\rightarrow\mathbb{C}$ of $(Y^{\mathrm{pol}},D^{\mathrm{pol}},L^{\mathrm{pol}})$, with central fiber given by a union of polarized toric surfaces $(\mathbb{P}^{2},\partial\mathbb{P}^{2},\mathcal{O}(1))$, and having $(P,\mathcal{P})$ as intersection complex. By appropriate scalings and refinements, one can ensure that after perturbing the decomposition $\widetilde{\mathcal{P}}$ on $\widetilde{P}$, all singularities still lie inside some edges. Then, as explained in [9], the momentum maps on the irreducible components $(\mathbb{P}^{2},\partial\mathbb{P}^{2},\mathcal{O}(1))$ of the central fiber glue to a fibration with base the intersection complex $(\widetilde{P},\widetilde{\mathcal{P}})$. By [9, Theorem 5.7], the almost toric fibration on $(Y^{\mathrm{pol}},D^{\mathrm{pol}},L^{\mathrm{pol}})$ can be obtained as a deformation of this fibration over $(\widetilde{P},\widetilde{\mathcal{P}})$.
@@ 5.3. Comparison with existing s-rules
The first proposed “s-rule” can be found in [18, §3.2]. If a decorated lattice polygon $(\overline{P},(L_{i})_{1\leq i\leq Q})$ satisfies the s-rule of [18, §3.2], then there exists in particular a polyhedral decomposition of $(\overline{P},(L_{i})_{1\leq i\leq Q})$ whose faces are either trapezoids or triangles whose three edges of the same lattice lengths. In particular, elementary triangles such as the one represented in the middle of Figure 4.6 are not allowed in the polyhedral decompositions considered in [18, §3.2]. It follows that the decorated polygon in Figure 7.10 does not satisfy the “s-rule” of [18, §3.2], whereas it admits a consistent regular polyhedral decomposition in the sense of Definition 5.4. The M-theory dual Calabi–Yau 3-fold to the corresponding web of $5$-branes with $7$-branes for this example is described in Example 7.9.
A modified version of the s-rule can be found in [94, §3.3]. In [94] it is proposed that a GTP defines a 5d SCFT if the s-rule in [94, §3.3] and the r-rule in [94, §3.4] are satisfied. Elementary triangles and elementary trapezoids are tiles in the sense of [94, Definition 5], and so it follows that a consistent decorated toric polygon $(P,(L_{i})_{1\leq i\leq Q})$ with a regular polyhedral decomposition $\overline{\mathcal{P}}$ satisfies the s-rule of [94, §3.3]. It also satisfies the r-rule $r\geq 0$ in [94, §3.4] since, by Lemma 5.8, the rank $r$ of the corresponding GTP is equal to $\mathrm{Area}(P)+1$, where $\mathrm{Area}(P)$ is the area of the associated Symington polygon. However, there are GTPs that satisfy both the s-rule and the r-rule but such that the corresponding decorated toric polygon is not consistent for any regular polyhedral decomposition. Moreover, even when the dual web is supersymmetric in the sense of Definition 4.1, the rank of the corresponding 5d SCFT is not necessarily equal to the rank $r$ of the GTP, for the general reason presented in Remark 4.9 and contrary to the expectation of [94] – see example below. It is an interesting open question to determine if the web dual to a GTP that satisfies both the s-and the r-rule of [94] is always supersymmetric in the sense of Definition 4.1.
@@
Consider the GTP $\overline{P}$ on the right of Figure 5.1. The polyhedral decomposition of $\overline{P}$ represented in Figure 5.1 satisfies the s-rule. The r-rule is also satisfied: the area of $\overline{P}$ is $\frac{37}{2}$ whereas the sum of squares of the parts of the partitions determined by the white dots is
@@
and so we have $r=\frac{37}{2}+1-\frac{37}{2}=1$ by [94, Equation 3.47]. Therefore, [94] predicts that the GTP $\overline{P}$ defines a rank one 5d SCFT. However, we argue that it is not the case as can be seen from the web picture. Pushing 7-branes as on the right of Figure 5.2 shows that the general curve $C\in|L|$ in the corresponding log Calabi–Yau surface with line bundle $(Y,D,L)$ is the disjoint union of a connected genus two curve and of a genus zero $(-2)$-curve. The arithmetic genus of this disconnected curve $C$ is indeed equal to $r=1$, but $C$ is not a connected genus one curve. Since $C$ is smooth, the web is supersymmetric according to Definition 4.1, and so we also expect the existence of a 5d SCFT. However, this 5d SCFT is not of rank one, but is obtained by adding one free hypermultiplet to a rank two SCFT, as explained following Definition 4.5. In fact, since a component of $C$ has negative self-intersection, it follows from Lemma 4.6 that the corresponding decorated toric polygon is not consistent – this can also be checked directly by inspecting the possible polyhedral decompositions.
Refer to caption2222444433332222333344442222
Figure 5.1. A non-consistent GTP that satisfies both the s-rule and the r-rule of [94].
Refer to caption4444333322223333444422222222
Figure 5.2. The general curve $C\in|L|$ is a disjoint union of a genus two curve and of a genus zero $(-2)$-curve.
@@ 6.1. Calabi-Yau 3-folds and mirror symmetry for log Calabi–Yau surfaces
    As in §4.1, the duality between Type IIB string theory on $S^{1}$ and M-theory on $T^{2}$ implies that the $S^{1}$-compactification of the 5d theory defined by $W^{\mathrm{asym}}$ in Type IIB string theory has a dual description as the $4d$ $\mathcal{N}=2$ theory defined by an M5-brane on $\mathbb{R}^{4}\times C^{\circ}$ in M-theory on $\mathbb{R}^{4}\times U\times\mathbb{R}^{3}$, where, denoting by $(Y,D,L)$ the log Calabi–Yau surface with line bundle associated to $W^{\mathrm{asym}}$ and $C$ a general curve in $|L|$, we have $U=Y\setminus D$ and $C^{\circ}=C\cap U$.
@@ 6.2. Calabi–Yau 3-folds as non-toric deformations of toric Calabi–Yau 3-folds
Let $W$ be a generic consistent web of 5-branes with 7-branes, defined by a web of 5-branes $\overline{W}$ and a 7-brane data. As explained in §3.4, $\overline{W}$ defines a degeneration $(\overline{\mathcal{Y}},\overline{\mathcal{D}},\overline{\mathcal{L}})\rightarrow\mathbb{C}$ of polarized log Calabi–Yau surfaces, and $W$ determines a degeneration $(\mathcal{Y},\mathcal{D},\mathcal{L})\rightarrow\mathbb{C}$ of log Calabi–Yau surfaces obtained from $(\overline{\mathcal{Y}},\overline{\mathcal{D}},\overline{\mathcal{L}})\rightarrow\mathbb{C}$ by blow-ups. By §4.3, while the line bundle $\mathcal{L}$ is not nef in general, it is possible to apply a finite sequence of M1-flop to obtain a degeneration $(\overline{\mathcal{Y}}^{\prime},\overline{\mathcal{D}}^{\prime},\overline{\mathcal{L}}^{\prime})\rightarrow\mathbb{C}$ of log Calabi–Yau surface with nef $\mathcal{L}^{\prime}$. By §4.6, this sequence of M1-flops is encoded into a modified web of 5-branes with 7-branes $W^{\prime}$ obtained from $W$ by “pushing fully” the 7-branes. As described in §5, the dual $W^{\prime}$ is a Symington polygon $P$ with polyhedral decomposition $\mathcal{P}$, obtained from the lattice polygon $\overline{P}$ with polyhedral decomposition $\overline{\mathcal{P}}$ dual to $\overline{W}$ by a series of cutting operations.
@@
Following [6] (see also [37]), we define below a Calabi–Yau 3-fold $\mathcal{X}$ associated to $W$, with a morphism $\pi:\mathcal{X}\rightarrow\Delta$ where $\Delta\subset\mathbb{C}$ is an analytic disk containing 00, such that the central fiber $\mathcal{X}_{0}:=\pi^{-1}(0)$ has dual intersection complex $(P,\mathcal{P})$, that is, intersection complex given by the polyhedral decomposition of $\mathbb{R}^{2}$ defined by the modified web $W^{\prime}$. We proceed in three steps described below: first, we construct for every vertex $v$ of $\mathcal{P}$ a log Calabi–Yau surface $(X_{0}^{v},\partial X_{0}^{v})$, then, we glue these log Calabi–Yau surfaces together to form a normal crossing surface $\mathcal{X}_{0}$, and finally, we construct a Calabi–Yau 3-fold $\mathcal{X}$ as total space of a one-parameter smoothing of $\mathcal{X}_{0}$.
[…]
@@
Figure 6.1. On the left, the star around a vertex $v$ of the polyhedral decomposition $\mathcal{P}$ of a Symington polygon $P$ depicted on the right.
@@
STEP I: Construction of the log Calabi–Yau surfaces $(X_{0}^{v},\partial X_{0}^{v})$. Let $v$ be a vertex of $\mathcal{P}$. If $v$ is not a singular point of the integral affine structure on the Symington polygon $\mathcal{P}$, then the star of $\mathcal{P}$ at $v$ can be viewed as a toric fan, and we denote by $(\overline{X}_{0}^{v},\partial\overline{X}_{0}^{v})$ the corresponding toric surface with its toric boundary divisor. If $v$ is a singular point of the integral affine structure on $\mathcal{P}$, then, the star of $\mathcal{P}$ at $v$ is obtained from the star of $v$ viewed as a vertex of $\overline{\mathcal{P}}$ by cutting out some cones, see Figure 6.1. Let $(\overline{X}_{0}^{v},\partial\overline{X}_{0}^{v})$ be the toric surface whose fan is the star of $\overline{\mathcal{P}}$ at $v$. Every cut cone $\mathcal{C}$ corresponds to a 0-dimensional stratum $x_{\mathcal{C}}$ of the toric boundary $\partial\overline{X}_{0}^{v}$. By construction of $(P,\mathcal{P})$ in §5.2, every cut cone $\mathcal{C}$ is a cone over an elementary triangle. Explicitly, if $O$ is the vertex of $\mathcal{C}$, and $A$, $B$ are the primitive integral points of the two rays of $\mathcal{C}$, then the triangle $OAB$ is an elementary triangle: the lattice length of the line segment $AB$ is equal to the lattice distance of $AB$ from $O$, see Figure 6.1. Every 0-dimensional stratum $x_{\mathcal{C}}$ in $(\overline{X}_{0}^{v},\partial\overline{X}_{0}^{v})$ admits a neighborhood isomorphic to the affine toric surface with $(\overline{X}_{\mathcal{C}},\partial{\overline{X}}_{\mathcal{C}})$ with cone $\mathcal{C}$. By [8, §7.3], $(\overline{X}_{\mathcal{C}},\partial{\overline{X}}_{\mathcal{C}})$ is a quotient singularity of the form
@@
where the weights of the $\mathbb{Z}/n^{2}\mathbb{Z}$-action are of the form $(1,an-1)$ with $\mathrm{gcd}(a,n)=1$. Moreover, the integer $n$ is the common lattice height and lattice base of the elementary triangle spanning the cone $\mathcal{C}$.
If $n=1$, then $(\overline{X}_{\mathcal{C}},\partial{\overline{X}}_{\mathcal{C}})$ is just $\mathbb{C}^{2}_{x,y}$ with its toric boundary $xy=0$. If $n>1$, then $(\overline{X}_{\mathcal{C}},\partial{\overline{X}}_{\mathcal{C}})$ is an elementary T-singularity in the terminology of [70, Proposition 3.10], also known as a Wahl singularity. Using the change of variables $x=u^{n}$, $y=v^{n}$, $z=uv$, this singularity can be equivalently described as the quotient of the hypersurface

### v2

@@ 4.6. Generic consistent webs and fully pushed 7-branes
Let $W$ be a generic consistent web of 5-branes with 7-branes. By Lemma 4.16, since $W$ is consistent, there exists a modified web of 5-branes with 7-branes $W^{\prime}$, obtained from $W$ by a finite sequence of pushing operation of 7-branes, such that the corresponding line bundle $\mathcal{L}^{\prime}$ on the degeneration of log Calabi–Yau surfaces $(\mathcal{Y}^{\prime},\mathcal{D}^{\prime},\mathcal{L}^{\prime})\rightarrow\mathbb{C}$ is nef. By the base point free theorem, there exists a morphism $c:(\mathcal{Y}^{\prime},\mathcal{D}^{\prime},\mathcal{L}^{\prime})\rightarrow((\mathcal{Y}^{\prime})^{\mathrm{pol}},(\mathcal{D}^{\prime})^{\mathrm{pol}},(\mathcal{L}^{\prime})^{\mathrm{pol}})$ contracting all curves $C^{\prime}$ in $\mathcal{Y}^{\prime}$ such that $\mathcal{L}^{\prime}\cdot C^{\prime}=0$, and so that the line bundle $(\mathcal{L}^{\prime})^{\mathrm{pol}}$ is ample. The polyhedral decomposition defined by the web $W^{\prime}$ is the dual intersection complex of the central fiber $((\mathcal{Y}^{\prime})^{\mathrm{pol}}_{0},(\mathcal{D}^{\prime})^{\mathrm{pol}}_{0},(\mathcal{L}^{\prime})^{\mathrm{pol}}_{0})$. Since $W^{\prime}$ is generic, the irreducible components of $((\mathcal{Y}^{\prime})^{\mathrm{pol}}_{0},(\mathcal{D}^{\prime})^{\mathrm{pol}}_{0},(\mathcal{L}^{\prime})^{\mathrm{pol}}_{0})$ are described by Lemma 4.27, and it follows that the 7-branes are disjoint from the web $W^{\prime}$, and that the vertices of $W^{\prime}$ are 3-valent of multiplicity one. To end the proof that the 7-branes are fully pushed, it remains to check condition ii) in Definition 4.28. If this condition were not satisfied, there would exist a $(p_{i},q_{i})$ 7-brane $x_{i}$ such that the half-line $x_{i}+\mathbb{R}_{<0}(p_{i},q_{i})$ does not intersect $W^{\prime}$. Then, there would exist an interior (1)(-1)-curve $E$ in a corner blow-up of $Y$ whose tropicalization is $x_{i}+\mathbb{R}_{<0}(p_{i},q_{i})$, and such that $L\cdot E=0$. But since $W$ is a consistent web, we have $L\cdot E>0$ by Definition 4.5 ii), which is a contradiction.
@@ 5.1. Consistent decorated toric polygons
As in §2.1.2, a polyhedral decomposition $\overline{\mathcal{P}}$ of a lattice polytope $\overline{P}$ is said to be regular if there exists a convex continuous piecewise linear function on $\overline{P}$ whose domains of linearity coincide with the faces of $\overline{\mathcal{P}}$.
@@
Let $(\overline{P},(L_{i})_{1\leq i\leq Q})$ be a decorated toric polygon equipped with a regular polyhedral decomposition $\overline{\mathcal{P}}$. We define a web of 5-branes with 7-branes $W$ to be dual to $(\overline{P},(L_{i})_{1\leq i\leq Q},\overline{\mathcal{P}})$ if there exists a convex continuous piecewise-linear function $\overline{\varphi}$ on $\overline{P}$ with domains of linearity $\overline{\mathcal{P}}$, such that, denoting by $\overline{W}$ the web of $5$-branes dual to $(\overline{\mathcal{P}},\overline{\varphi})$, $W$ is obtained from $\overline{W}$ by adding for every $1\leq i\leq Q$ a 7-brane ending the leg of $\overline{W}$ corresponding to $L_{i}$.
Recall from §3.3 that any web of $5$-branes with $7$-branes determines a log Calabi surface with line bundle, denoted $(Y,D,L)$. In what follows, the expression the log Calabi surface with line bundle associated to a decorated toric polygon with a regular polyhedral decomposition refers to the corresponding data $(Y,D,L)$ associated with the dual web.
@@
    For every $1\leq i\leq Q$, a set of faces $\mathcal{F}_{\Psi,i}\subseteq\overline{\mathcal{P}}^{[2]}$. We refer to a face in $\mathcal{F}_{\Psi,i}$ as a $(\Psi,i)$-colored face”, and require that $\mathcal{F}_{\Psi,i}\cap\mathcal{F}_{\Psi,j}=\emptyset$, if $i\neq j$.
@@
    For every $1\leq i\leq Q$, an element $A_{\Psi,i}\in\overline{\mathcal{P}}^{[0]}\cup\overline{\mathcal{P}}^{[1]}$, referred to as the $(\Psi,i)$-active cell”.
@@
    $\mathcal{F}_{\Psi_{\mathrm{in}},i}$$\emptyset$, and;
@@
    for all $1\leq j\leq k-1$, $\sigma_{j}$ is an elementary trapezoid, with two parallel edges $e_{in,j}$ and $e_{out,j}$, which were once active cells, and with two other edges $e_{l,j}$ and $e_{r,j}$ such that $e_{l,j}\sim_{\Psi}e_{r,j}$.
@@
Moreover, we have $|e_{in,1}|=|L_{i}|$, and, for every $1\leq j\leq k-1$, we have $|e_{out,j}|=|e_{in,j+1}|$.
@@
Example 5.8.
Figure 18 depicts a consistent decorated toric polygon with a regular polyhedral decomposition, along with a final state. This final state is obtained by first applying a trapezoid-coloring move, and then a triangle-coloring move. The trapezoid-coloring move is dual to the 7-brane crossing a 4-valent vertex of the web, whereas the triangle-coloring move is dual to the 7-brane crossing a 3-valent vertex of the web.
Figure 18. On the left, a consistent decorated toric polygon with a regular polyhedral decomposition and its dual web. On the right, a corresponding coloring and the dual web with fully pushed 7-branes.
@@ 5.2. Symington polygons
[…]
@@ 5.3. Comparison with existing s-rules
The first proposed “s-rule” can be found in [21, §3.2]. If a decorated lattice polygon $(\overline{P},(L_{i})_{1\leq i\leq Q})$ satisfies the s-rule of [21, §3.2], then there exists in particular a polyhedral decomposition of $(\overline{P},(L_{i})_{1\leq i\leq Q})$ whose faces are either trapezoids or triangles whose three edges of the same lattice lengths. In particular, elementary triangles such as the one represented in the middle of Figure 13 are not allowed in the polyhedral decompositions considered in [21, §3.2]. It follows that the decorated polygon in Figure 37 does not satisfy the “s-rule” of [21, §3.2], whereas it admits a consistent regular polyhedral decomposition in the sense of Definition 5.4. The M-theory dual Calabi–Yau 3-fold to the corresponding web of $5$-branes with $7$-branes for this example is described in Example 7.9.
A modified version of the s-rule can be found in [97, §3.3]. In [97] it is proposed that a GTP defines a 5d SCFT if the s-rule in [97, §3.3] and the r-rule in [97, §3.4] are satisfied. Elementary triangles and elementary trapezoids are tiles in the sense of [97, Definition 5], and so it follows that a consistent decorated toric polygon $(P,(L_{i})_{1\leq i\leq Q})$ with a regular polyhedral decomposition $\overline{\mathcal{P}}$ satisfies the s-rule of [97, §3.3]. It also satisfies the r-rule $r\geq 0$ in [97, §3.4] since, by Lemma 5.9, the rank $r$ of the corresponding GTP is equal to $\mathrm{Area}(P)+1$, where $\mathrm{Area}(P)$ is the area of the associated Symington polygon. However, there are GTPs that satisfy both the s-rule and the r-rule but such that the corresponding decorated toric polygon is not consistent for any regular polyhedral decomposition. Moreover, even when the dual web is supersymmetric in the sense of Definition 4.1, the rank of the corresponding 5d SCFT is not necessarily equal to the rank $r$ of the GTP, for the general reason presented in Remark 4.9 and contrary to the expectation of [97] – see example below. It is an interesting open question to determine if the web dual to a GTP that satisfies both the s-and the r-rule of [97] is always supersymmetric in the sense of Definition 4.1.
@@
Consider the GTP $\overline{P}$ on the right of Figure 19. The polyhedral decomposition of $\overline{P}$ represented in Figure 19 satisfies the s-rule. The r-rule is also satisfied: the area of $\overline{P}$ is $\frac{37}{2}$ whereas the sum of squares of the parts of the partitions determined by the white dots is
@@
and so we have $r=\frac{37}{2}+1-\frac{37}{2}=1$ by [97, Equation 3.47]. Therefore, [97] predicts that the GTP $\overline{P}$ defines a rank one 5d SCFT. However, we argue that it is not the case as can be seen from the web picture. Pushing 7-branes as on the right of Figure 20 shows that the general curve $C\in|L|$ in the corresponding log Calabi–Yau surface with line bundle $(Y,D,L)$ is the disjoint union of a connected genus two curve and of a genus zero (2)(-2)-curve. The arithmetic genus of this disconnected curve $C$ is indeed equal to $r=1$, but $C$ is not a connected genus one curve. Since $C$ is smooth, the web is supersymmetric according to Definition 4.1, and so we also expect the existence of a 5d SCFT. However, this 5d SCFT is not of rank one, but is obtained by adding one free hypermultiplet to a rank two SCFT, as explained following Definition 4.5. In fact, since a component of $C$ has negative self-intersection, it follows from Lemma 4.6 that the corresponding decorated toric polygon is not consistent – this can also be checked directly by inspecting the possible polyhedral decompositions.
Figure 19. A non-consistent GTP that satisfies both the s-rule and the r-rule of [97].
Figure 20. The general curve $C\in|L|$ is a disjoint union of a genus two curve and of a genus zero (2)(-2)-curve.
@@ 6.1. Calabi-Yau 3-folds and mirror symmetry for log Calabi–Yau surfaces
    Denote by $(Y,D,L)$ the log Calabi–Yau surface with line bundle associated to $W^{\mathrm{asym}}$, $C$ a general curve in $|L|$, $U=Y\setminus D$ and $C^{\circ}=C\cap U$. Then, as discussed in §4.1, the duality between Type IIB string theory on $S^{1}$ and M-theory on $T^{2}$ implies that the $S^{1}$-compactification of the 5d theory defined by $W^{\mathrm{asym}}$ in Type IIB string theory has a dual description as the $4d$ $\mathcal{N}=2$ theory defined by an M5-brane on $\mathbb{R}^{4}\times C^{\circ}$ in M-theory on $\mathbb{R}^{4}\times U\times\mathbb{R}^{3}$.
@@ 6.2. Calabi–Yau 3-folds as non-toric deformations of toric Calabi–Yau 3-folds
Let $W$ be a generic consistent web of 5-branes with 7-branes, defined by a web of 5-branes $\overline{W}$ and a 7-brane data. As explained in §3.4, $\overline{W}$ defines a degeneration $(\overline{\mathcal{Y}},\overline{\mathcal{D}},\overline{\mathcal{L}})\rightarrow\mathbb{C}$ of polarized log Calabi–Yau surfaces, and $W$ determines a degeneration $(\mathcal{Y},\mathcal{D},\mathcal{L})\rightarrow\mathbb{C}$ of log Calabi–Yau surfaces obtained from $(\overline{\mathcal{Y}},\overline{\mathcal{D}},\overline{\mathcal{L}})\rightarrow\mathbb{C}$ by blow-ups. By §4.3, although the line bundle $\mathcal{L}$ is generally not nef, one can perform a finite sequence of M1-flop to produce a degeneration $(\overline{\mathcal{Y}}^{\prime},\overline{\mathcal{D}}^{\prime},\overline{\mathcal{L}}^{\prime})\rightarrow\mathbb{C}$ of log Calabi–Yau surface with nef $\mathcal{L}^{\prime}$. According to §4.6, this sequence of M1-flops is encoded into a modified web of 5-branes with 7-branes $W^{\prime}$ obtained from $W$ by “fully pushing” the 7-branes. As described in §5, the dual $W^{\prime}$ is a Symington polygon $P$ with polyhedral decomposition $\mathcal{P}$, obtained from the lattice polygon $\overline{P}$ with polyhedral decomposition $\overline{\mathcal{P}}$ dual to $\overline{W}$ by a series of cutting operations.
@@
Following [6] (see also [40]), we define below a Calabi–Yau 3-fold $\mathcal{X}$ associated to $W$, with a morphism $\pi:\mathcal{X}\rightarrow\Delta$ where $\Delta\subset\mathbb{C}$ is an analytic disk containing 0, such that the central fiber $\mathcal{X}_{0}:=\pi^{-1}(0)$ has dual intersection complex $(P,\mathcal{P})$, that is, intersection complex given by the polyhedral decomposition of $\mathbb{R}^{2}$ defined by the modified web $W^{\prime}$. We proceed in three steps described below. First, for each vertex $v$ of $\mathcal{P}$, we construct a log Calabi–Yau surface $(X_{0}^{v},\partial X_{0}^{v})$. Next, we glue these log Calabi–Yau surfaces together to form a normal crossing surface $\mathcal{X}_{0}$. Finally, we construct a Calabi–Yau 3-fold $\mathcal{X}$ as the total space of a one-parameter smoothing of $\mathcal{X}_{0}$.
@@
Figure 21. On the left, the star around a vertex $v$ of the polyhedral decomposition $\mathcal{P}$ of a Symington polygon $P$ depicted on the right.
@@
STEP I: Construction of the log Calabi–Yau surfaces $(X_{0}^{v},\partial X_{0}^{v})$. Let $v$ be a vertex of $\mathcal{P}$. If $v$ is not a singular point of the integral affine structure on the Symington polygon $\mathcal{P}$, then the star of $\mathcal{P}$ at $v$ can be viewed as a toric fan, and we denote by $(\overline{X}_{0}^{v},\partial\overline{X}_{0}^{v})$ the corresponding toric surface with its toric boundary divisor. If $v$ is a singular point of the integral affine structure on $\mathcal{P}$, then, the star of $\mathcal{P}$ at $v$ is obtained from the star of $v$ viewed as a vertex of $\overline{\mathcal{P}}$ by cutting out some cones, see Figure 21. Let $(\overline{X}_{0}^{v},\partial\overline{X}_{0}^{v})$ be the toric surface whose fan is the star of $\overline{\mathcal{P}}$ at $v$. Every cut cone $\mathcal{C}$ corresponds to a 0-dimensional stratum $x_{\mathcal{C}}$ of the toric boundary $\partial\overline{X}_{0}^{v}$. By construction of $(P,\mathcal{P})$ in §5.2, every cut cone $\mathcal{C}$ is a cone over an elementary triangle. Explicitly, if $O$ is the vertex of $\mathcal{C}$, and $A$, $B$ are the primitive integral points of the two rays of $\mathcal{C}$, then the triangle $OAB$ is an elementary triangle: the lattice length of the line segment $AB$ is equal to the lattice distance of $AB$ from $O$, see Figure 21. Every 0-dimensional stratum $x_{\mathcal{C}}$ in $(\overline{X}_{0}^{v},\partial\overline{X}_{0}^{v})$ admits a neighborhood isomorphic to the affine toric surface with $(\overline{X}_{\mathcal{C}},\partial{\overline{X}}_{\mathcal{C}})$ with cone $\mathcal{C}$. By [8, §7.3], $(\overline{X}_{\mathcal{C}},\partial{\overline{X}}_{\mathcal{C}})$ is a quotient singularity of the form
@@
where the weights of the $\mathbb{Z}/n^{2}\mathbb{Z}$-action are of the form (1,an1)(1,an-1) with $\mathrm{gcd}(a,n)=1$. Moreover, the integer $n$ is the common lattice height and lattice base of the elementary triangle spanning the cone $\mathcal{C}$.
If $n=1$, then $(\overline{X}_{\mathcal{C}},\partial{\overline{X}}_{\mathcal{C}})$ is simply $\mathbb{C}^{2}_{x,y}$ with its toric boundary $xy=0$. If $n>1$, then $(\overline{X}_{\mathcal{C}},\partial{\overline{X}}_{\mathcal{C}})$ is an elementary T-singularity in the terminology of [73, Proposition 3.10], also known as a Wahl singularity. Using the change of variables $x=u^{n}$, $y=v^{n}$, $z=uv$, this singularity can be equivalently described as the quotient of the hypersurface